OpenStudy (anonymous):

Input the equation for the following line in standard form Ax+By=C Reduce all fractional answers to lowest terms. A line that passes through the origin and is parallel to x+y=6

OpenStudy (anonymous):

When you know that a line is parallel to another line, you can automatically say that their slopes are the same. It is easy to see the slope for x+y=6 by changing it to y=-x+6 by subtracting x to the other side. Also, you know that the line passes through the origin, so you can substitute x for 0 and y for 0 in Ax+By=C so A(0)+B(0)=C and we know that anything multiplied by 0 is 0 so 0=C or C=0 With the slope and y-intercept, you can create your equation.